Five black balls numbered 1,2,3,4, and 5 and seven white balls numbered 1,2,3,4,5,6, and 7 are placed in an urn. If one is chosen at random, what is the probability that it is numbered 5 or that it is white? and What is the probability that it is numbered 5 given that it is white?
Answer by stanbon(75887)

Five black balls numbered 1,2,3,4, and 5 and seven white balls numbered 1,2,3,4,5,6, and 7 are placed in an urn.
------------------------------------------
If one is chosen at random, what is the probability that it is numbered 5 or that it is white?
Prob(5 or white) = Prob(5) + Prob(white) - Prob(5 and white)
=2/12 + 7/12 - 1/12 = 8/12= 2/3
------------------------------------------
What is the probability that it is numbered 5 given that it is white?
Prob(5 given white) = [Prob(5 and white)]/[Prob(white)]
=[1/12]/[7/12] = 1/7
--------------------------------------
Cheers,
Stan H.